Nichrome wire of cross-sectional radius 0.798 mm is to be used inwinding a heating coil. If the coil must carry a current of 8.73 A when a voltage of 1.3 x 102 V is applied across its ends, find:

a) the required resistance of the coil and
121.27Ω

147.46Ω

138.73Ω

11349Ω
b) the length of wire you must use to wind the coil.

31.78 m

7.94m

19.86m

51.64m
Answer #1

r = 0.798 mm = 7.98\times 10^{-4} m

I = 8.73 A

AV = 1.3 x 102V

a)

the resistance R = \frac{\bigtriangleup V }{I}

R = \frac{1.3\times 10^{2}V }{8.73 A}

R = 14.89 \Omega

b)

the resistance R = \frac{\rho L }{A}

L = \frac{RA }{\rho }

A = \pi r^{2}

L = \frac{R\left ( \pi r^{2} \right ) }{\rho }

L = \frac{14.89\Omega \times \left ( \pi \left ( 7.98\times 10^{-4}m \right )^{2} \right ) }{150\times 10^{-8}\Omega .m }

L = 19.86 m